At XYZ University, another group of students was selected and asked how much of their weekly allowance they spent in buying mobile phone load. The following is the list of amounts spent: Php 200, 180, 30, 20, 10, 160, 150, 80. find the standard deviation



1
Expert's answer
2021-09-27T15:02:21-0400

Mean: "\frac{1}{n}\Sigma x_i=\frac{830}{8}=103.75."

Variance: "\sigma^2=\frac{1}{n} \Sigma(x_i-\mu)^2=\frac{42187.5}{8}=5273.4375."

Standard deviation: "\sigma=\sqrt{5273.4375}=72.6184."